[The following information applies to the questions displayed below.]

Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 14,250 $ 17,550 $ 31,800
Estimated variable manufacturing overhead per machine-hour $ 3.10 $ 3.90
Job P Job Q
Direct materials $ 30,000 $ 16,500
Direct labor cost $ 34,600 $ 14,300
Actual machine-hours used:
Molding 3,400 2,500
Fabrication 2,300 2,600
Total 5,700 5,100

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

Foundational 2-12

12. If Job P included 20 units, what was its unit product cost?

If Job Q included 30 units, what was its unit product cost?

Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q?

What was Sweeten Company’s cost of goods sold for March?
Answer #1
Overhead rates :
Molding 8.80 =3.1+(14250/2500)
Fabrication 15.60 =3.9+(17550/1500)
Job P Job Q
Direct materials 30000 16500
Direct labor cost 34600 14300
Manufacturing overhead applied:
Molding 29920 22000
Fabrication 35880 40560
Total manufacturing cost 130400 93360
12
Total manufacturing cost Job P 130400
Divide by units 20
Unit product cost Job P 6520
13
Total manufacturing cost Job Q 93360
Divide by units 30
Unit product cost Job Q 3112
14
Job P Job Q
Total manufacturing cost 130400 93360
Add: Markup @ 80% 104320 74688
Total price for the job 234720 168048
Divide by units 20 30
Selling price per unit 11736 5602
15
Cost of goods sold for March 223760 =130400+93360